Blades Inc., US firm has agreed to sell its product to Regency, EU firm and payment is Euro 2 million. Blades Inc has agreement that the amount will be paid in 30 days. Blades Inc wants to avoid the FX risk by using the US and EU money market. Following information is available to Blades Inc.

Current sport rate = $1.15/euroEU borrowing rate = 8% annual rateUS investing rate = 6% annual rate

Explain how Blades Inc can conduct money market hedge? A. Borrow euro 1,986,689.18 from a Bank in Europe B. Borrow euro 1,990,050 from a Bank in US C. Borrow euro 1,851,851.85 from a Bank in Europe D. Borrow euro 1,886,792.45 from a Bank in US
Answer #1

Step 1:Since we are having a Receivable of 2 Million Euros after 1 month, So inorder to avoid the Foreign Exchange Risk we have to borrow the Present value of 2 Million Euros

EU Borrowing Rate = 8%

EU Borrowing rate for 1 Month = 8% /12 = 0.67%

US Investing Rate = 6%

US Investing rate for 1 Month = 6% /12 = 0.5% Per month

We have to borrow the Present value of 2 Mmillion Euro @ 0.6667% per month and invest in US at 0.5% per month

Present value of 2 Miillion EURO s = 2 Million EURO/ ( 1+0.67%)

= 2 Milion EURO / 1.0067

= 1.9866891 Million EURO

Present value of 2 Million EURO = 19,86689.91 EURO

So we have to borrow 19,86,689.91 EURO

So Option a ) Borrow 19,86,689.18 is the Correct answer.

Step 2:

Convert 19,86,689.91 into USD

1 EURO = 1.15 USD

19,86689.91 EURO = 1.15*19,86689.91 USD

= 22,84693.3965 USD

Hence we can get 22,84,693.3965 USD

Step 3:

We can invest 22,84693.3965 USD at 0.5% for 1 month.

After 1 month then the investment will be = 22,84693.3965*1.005

= 2296116.86 USD.

Step 4:

Loan balance after 1 Month = 19,86,689.91*1.0067

= 20,00000 EURO

After 1 month we have a receivable of 2 Million EURO, So we can repay the loan amount.

Note : Answer ends with step 1

If you are having any doubt,please post a Comment.

Thank you.Please rate it.